# Planck stars, White Holes, Remnants and Planck-mass quasi-particles The quantum gravity phase in black holes’ evolution and its manifestations

### v3

@@ IV.2 Direct detection
A local dark matter density of the order of $0.01M_{\odot}/pc^{3}$ corresponds to approximately one Planck-scale white hole per each $10.000Km^{3}$. These objects are presumably moving fast with respect to our local frame, since we are rotating with the galaxy at hundreds of Km per second, while dark matter probably isn’t. This gives a very rough estimates of a few Planck scale particles per $m^{2}$ per year flying by us. Can we detect them?
If the sole interaction of remnants is gravitational, they are very good dark matter candidates, but for this same reason direct detection using classical sensing is challenging Carney2018a , due to the extreme weakness of the gravitational interaction. However, there may be a quantum technology that could open a window to do the detection. In fact, recent developments in the area of table-top experiments involving gravity and quantum phenomena (see for instance Christodoulou2022 for up do date references) open the theoretical possibility of direct detection of purely-gravitationally-interacting dark matter particles. A idealized detector where the center of detector mass is set in a superposition of locations and a more concrete tentative protocol, which employs Josephson junctions, have been illustrated Perez2023 .
@@ IV.3 Cosmological implications
Remnants could be produced by primordial black holes formed in the early universe. The lower bound of the lifetime of the remnants is of the order or lower to $m_{o}^{4}$. If we assume that lifetime is precisely of the order of the lower bound $m_{o}^{4}$ (as mentioned, this may not be correct), then for these objects to be still present now we need their lifetime to be larger or equal than the Hubble time $T_{H}$, that is
@@ IV.4 Erebons
The possibility that black holes could live across the big bounce and represent a component of dark matter has been considered in Carr2017 . Roger Penrose has coined the name erebons, from the Greek god of darkness Erebos, to refer to matter crossing over from one eon to the successive one Penrose2017 in his cyclic cosmology Penrose2012 . Large black holes evaporated before the bounce could have given rise to a population of Planck-size white holes remnants that has crossed the bounce and formed what we see today as dark matter. For this to happen, their density should have been sufficient to balance the huge dilution in an eventual inflationary phase.
An intriguing aspect of this scenario is the speculative idea that it might addresses the apparent low-entropy of the initial cosmological state Rovelli2018i . This can be consistent and can represent a concrete realisation of the perspectival interpretation of entropy suggested in Rovelli:2015 . If the cosmos at the big bounce was finely dotted by white holes with large interiors, then the gravitational field was not in the very improbable low entropy homogeneous or nearly homogeneous configuration. It was in a high entropy crumpled configuration. But being outside all white holes, we are in a special place, and from the special perspective of this place we see the universe under a coarse grain which defines an entropy that was low in the past.
@@ IV.5 Modeling remnants emission
Classically, this a very low power emission. Quantum mechanics, it is given by a weak decay probability, akin radioactivity. This process was tentatively modelled in Kazemian2022a , which we summarize here.
@@
Here $A$ is the area of the horizon at formation. Information is emitted in form of radiation. Since the radiation is emitted radially, it can be modeled as a uniform one-dimensional gas of photons. Assume for simplicity that the radiation emitted emitted by the surface of the remnant during the lifetime $\tau$, is in thermal equilibrium Marolf2017a . At the end of the remnant lifetime the radiation covers a length $L=\tau$. The energy $E$ available for this gas is only that of the mass of the remnant, which is of the order of the Planck mass, namely unit in natural units.
@@
The life-time of a white hole would be equal to time required for the photons to travel a distant $L$. We therefore have
@@
A population of black holes formed at a time $t=0$, with mass $m$ and uniformly distributed in space that evaporate around time $\tau_{B}\sim m^{3}$ as predicted by Hawking radiation theory, and survive as white hole remnants for a time $\tau_{W}$ as in (56) emits a steady radiation between times $\tau_{B}$ and $\tau_{B}+\tau_{W}$. For $\tau_{B}<t<\tau_{B}+\tau_{W}$ an observer would observe the radiation density would evolve in time as
@@
Quantum mechanics alters this picture a bit. A continuous energy emission as the one described above would imply a continuously decreases of the white hole horizon area, below the Planck area. But this is not permitted by LQG, because area is quantized.
Rather, a remnant with near-Planckian mass and area must make a single quantum leap into radiation, as in conventional nuclear radioactivity, where emission is realised by individual quantized quanta governed by a probability distribution Martin2019 .
@@
The number of photons emitted by a single remnant is given in (55). This conclusion is relevant in view of an old objection to the remnant scenario, that contributed to its abandonment in the Nineties. The objection was that the large number of remnant internal states would make them too easy to produce in particle physics experiments. Here we see why that conclusion was to quick. The effective vertex responsible for a remnant production would be
@@
In the cosmological standard model, primordial black holes may have formed at reheating. To get a sense of the characteristic of the diffuse radiation remnants may emit, one can estimate its parameter in this simplest case Kazemian2022a . This model is entirely determined by a single parameter or order of unity, that can be taken to be
@@
Other cosmological scenarios, such as big bounce or matter bounce scenarios, remnants might also –and perhaps better– accommodate remnants as a component of dark matter.

### v4

@@ IV.2 Direct detection
A local dark matter density of the order of $0.01M_{\odot}/pc^{3}$ corresponds to approximately one Planck-scale white hole per each $10.000Km^{3}$. These objects are presumably moving fast with respect to our local frame, since we are rotating with the galaxy at hundreds of Km per second, while dark matter models suggest that it isn’t. This gives a very rough estimates of a few Planck scale particles per $m^{2}$ per year flying by us. Can we detect them?
If the sole interaction of remnants is gravitational, they are very good dark matter candidates, but for this same reason direct detection using classical sensing is challenging Carney2018a , due to the extreme weakness of the gravitational interaction. However, there may be a quantum technology that could open a window to do the detection. In fact, recent developments in the area of table-top experiments involving gravity and quantum phenomena (see for instance Christodoulou2022 for up-do-date references) open the theoretical possibility of direct detection of purely-gravitationally-interacting dark matter particles. An idealized detector where the center of detector mass is set in a superposition of locations and a more concrete tentative protocol, which employs Josephson junctions, have been illustrated Perez2023 .
@@ IV.3 Cosmological implications
Remnants could be produced by primordial black holes formed in the early universe. The lower bound of the lifetime of the remnants is of the order of or lower than $m_{o}^{4}$. If we assume that lifetime is precisely of the order of the lower bound $m_{o}^{4}$ (as mentioned, this may not be correct), then for these objects to still be present now we need their lifetime to be larger or equal than the Hubble time $T_{H}$, that is
@@ IV.4 Erebons
The possibility that black holes could live across the big bounce and represent a component of dark matter has been considered in Carr2017 . Roger Penrose has coined the name erebons, from the Greek god of darkness Erebos, to refer to matter crossing over from one eon to the successive one Penrose2017 in his cyclic cosmology Penrose2012 . Large black holes evaporated before the bounce could have given rise to a population of Planck-size white hole remnants that has crossed the bounce and formed what we see today as dark matter. For this to happen, their density should have been sufficient to balance the huge dilution in an eventual inflationary phase.
An intriguing aspect of this scenario is the speculative idea that it might address the apparent low-entropy of the initial cosmological state Rovelli2018i . This can be consistent and can represent a concrete realization of the perspectival interpretation of entropy suggested in Rovelli:2015 . If the cosmos at the big bounce was finely dotted by white holes with large interiors, then the gravitational field was not in the very improbable low entropy homogeneous or nearly homogeneous configuration. It was in a high entropy crumpled configuration. But being outside all white holes, we are in a special place, and from the special perspective of this place we see the universe under a coarse grain which defines an entropy that was low in the past.
@@ IV.5 Modeling remnants emission
Classically, this a very low power emission. Quantum mechanically, it is given by a weak decay probability, akin to radioactivity. This process was tentatively modelled in Kazemian2022a , which we summarize here.
@@
Here $A$ is the area of the horizon at formation. Information is emitted in the form of radiation. Since the radiation is emitted radially, it can be modelled as a uniform one-dimensional gas of photons. Assume for simplicity that the radiation emitted by the surface of the remnant during the lifetime $\tau$, is in thermal equilibrium Marolf2017a . At the end of the remnant lifetime the radiation covers a length $L=\tau$. The energy $E$ available for this gas is only that of the mass of the remnant, which is of the order of the Planck mass, namely unity in natural units.
@@
The lifetime of a white hole would be equal to the time required for the photons to travel a distant $L$. We therefore have
@@
A population of black holes formed at a time $t=0$, with mass $m$ and uniformly distributed in space that evaporate around time $\tau_{B}\sim m^{3}$ as predicted by Hawking radiation theory, and survive as white hole remnants for a time $\tau_{W}$ as in (55) emits a steady radiation between times $\tau_{B}$ and $\tau_{B}+\tau_{W}$. For $\tau_{B}<t<\tau_{B}+\tau_{W}$ an observer would observe that the radiation density would evolve in time as
@@
Quantum mechanics alters this picture a bit. A continuous energy emission as the one described above would imply a continuous decrease of the white hole horizon area, below the Planck area. But this is not permitted by LQG, because area is quantized.
Rather, a remnant with near-Planckian mass and area must make a single quantum leap into radiation, as in conventional nuclear radioactivity, where emission is realized by individual quantized quanta governed by a probability distribution Martin2019 .
@@
The number of photons emitted by a single remnant is given in (54). This conclusion is relevant in view of an old objection to the remnant scenario, that contributed to its abandonment in the Nineties. The objection was that the large number of remnant internal states would make them too easy to produce in particle physics experiments. Here we see why that conclusion was too quick. The effective vertex responsible for a remnant production would be
@@
In the cosmological standard model, primordial black holes may have formed at reheating. To get a sense of the characteristic of the diffuse radiation that remnants may emit, one can estimate its parameter in the simplest case Kazemian2022a . This model is entirely determined by a single parameter of order of unity, that can be taken to be
@@
In other cosmological scenarios, such as big bounce or matter bounce scenarios, remnants might also –and perhaps better– accommodate remnants as a component of dark matter.
V Acknowledgments
We thank Richard Werthamer for his accurate corrections to the manuscript. This work was made possible through the support of the ID# 62312 grant from the John Templeton Foundation, as part of the project “The Quantum Information Structure of Spacetime” (QISS). CR acknowledges support from the Perimeter Institute for Theoretical Physics through its distinguished research chair program. FV acknowledge support from the Perimeter Institute for Theoretical Physics through its affiliation program. Research at Perimeter Institute is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Economic Development and Innovation. Research in FV’s research group at Western University is supported by the Canada Research Chairs Program and by the Natural Science and Engineering Council of Canada (NSERC) through the Discovery Grant “Loop Quantum Gravity: from Computation to Phenomenology.” We acknowledge the Anishinaabek, Haudenosaunee, Lūnaapéewak, Attawandaron, and neutral peoples, on whose traditional lands Western University and the Perimeter Institute are located.